A curve is described by 𝑦=𝑝π‘₯2+π‘žπ‘₯, where 𝑝 and π‘ž are constants.

a. Find an expression for the gradient of this curve at any point.

b. Given that at the point (1,βˆ’2) the gradient is 6, calculate the values of 𝑝 and π‘ž.

c. Show that the equation of the normal to the curve at the point (1,βˆ’2) can be written as π‘₯+6𝑦+11=0.


1
Expert's answer
2022-05-31T12:47:38-0400

a.

"y=px^2+qx"

"grad (y)=(px^2+qx)'=2px+q"

b.


"-2=p(1)^2+q(1)""2p(1)+q=6"

"q=-2-p""2p-2-p=6"

"p=8, q=-10"

c.

Find the slope of the normal at the point "(1, -2)"


"slope=m=-\dfrac{1}{grad (y)}=-\dfrac{1}{6}"

The equation of the normal at the point "(1,-2)" is


"y-(-2)=-\dfrac{1}{6}(x-1)"

"6y+12=-x+1"

Then he equation of the normal to the curve at the point "(1,\u22122)" can be written as


"x+6y+11=0"